# EventFlow: Forecasting Temporal Point Processes with Flow Matching

### v2

@@ E.1 MARE
Diffusion $1.95{\scriptstyle\pm 0.48}$ $1.28{\scriptstyle\pm 0.09}$ $1.12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
@@ E.2 MMDs
NHP $7.2{\scriptstyle\pm 0.2}$ $2.2{\scriptstyle\pm 1.6}$ $22.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
@@
Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $22.5{\scriptstyle\pm 0.5}$ $2.2{\scriptstyle\pm 0.8}$
@@ E.3 MSEs
IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $1.24{\scriptstyle\pm 0.16}$ $1.11{\scriptstyle\pm 0.17}$
NHP $0.89{\scriptstyle\pm 0.09}$ $0.53{\scriptstyle\pm 0.24}$ $\mathbf{0.0022}{\scriptstyle\pm 0.0007}$ $0.31{\scriptstyle\pm 0.12}$ $2.00{\scriptstyle\pm 0.30}$ $1.30{\scriptstyle\pm 0.26}$ $\underline{1.03}{\scriptstyle\pm 0.35}$
Diffusion $\mathbf{0.61}{\scriptstyle\pm 0.10}$ $0.33{\scriptstyle\pm 0.04}$ $\underline{0.0037}{\scriptstyle\pm 0.0012}$ $0.23{\scriptstyle\pm 0.14}$ $\mathbf{1.30}{\scriptstyle\pm 0.21}$ $\mathbf{0.86}{\scriptstyle\pm 0.18}$ $\mathbf{0.92}{\scriptstyle\pm 0.20}$
@@
EventFlow (10 NFEs) ${0.60}{\scriptstyle\pm 0.02}$ $1.01{\scriptstyle\pm 0.31}$ $0.0113{\scriptstyle\pm 0.0016}$ ${0.16}{\scriptstyle\pm 0.01}$ $1.93{\scriptstyle\pm 0.07}$ ${0.77}{\scriptstyle\pm 0.05}$ ${1.14}{\scriptstyle\pm 0.07}$

### v3

@@ E.1 MARE
Diffusion $1.95{\scriptstyle\pm 0.48}$ $28{\scriptstyle\pm 0.09} 1.1$$12{\scriptstyle\pm 0.56}$ $0.49{\scriptstyle\pm 0.07}$ $0.66{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.07}$ $0.72{\scriptstyle\pm 0.07}$
@@ E.2 MMDs
NHP $7.2{\scriptstyle\pm 0.2}$ $2{\scriptstyle\pm 1.6} 22.$$2.5{\scriptstyle\pm 0.3}$ $\underline{5.0}{\scriptstyle\pm 0.1}$ $7.3{\scriptstyle\pm 0.7}$ $6.7{\scriptstyle\pm 1.5}$ $6.1{\scriptstyle\pm 2.3}$
@@
Add-and-Thin $\underline{1.9}{\scriptstyle\pm 0.5}$ $2.5{\scriptstyle\pm 0.3}$ $\underline{2.6}{\scriptstyle\pm 0.5}$ $7.4{\scriptstyle\pm 1.2}$ $2.5{\scriptstyle\pm 0.5} 2.2$$2{\scriptstyle\pm 0.8}$
@@ E.3 MSEs
IFTPP $0.85{\scriptstyle\pm 0.05}$ $\underline{0.32}{\scriptstyle\pm 0.03}$ $0.0047{\scriptstyle\pm 0.0006}$ $0.22{\scriptstyle\pm 0.03}$ $1.74{\scriptstyle\pm 0.10}$ $24{\scriptstyle\pm 0.16} 1.1$$11{\scriptstyle\pm 0.17}$
[…]
@@
[…]
@@ E.4 Forecasting TPPs
E.5 Runtimes
In Table 21, we report the wall-clock time required to generate 1 0001\,000 sequences for each method. As expected, NHP and IFTPP achieve the fastest generation times, though this comes at the cost of substantially lower forecast quality (see Figure 2). Both Diffusion and Add-and-Thin are relatively slower due to their iterative refinement procedures. EventFlow (with 1 NFE) attains state-of-the-art forecasting accuracy while maintaining generation times comparable to IFTPP on datasets with moderate sequence lengths (Yelp-A, Yelp-M, Taxi). On datasets with longer sequences (e.g., Reddit-S), the transformer-based architecture of EventFlow naturally incurs higher computational cost. Although EventFlow with 25 NFEs is slower, Table 1 shows that a single NFE already suffices to reach SOTA performance. We note that these runtime comparisons are meant as approximate indicators, since implementation and architectural details can substantially influence wall-clock times. Overall, EventFlow achieves competitive generation efficiency despite not being explicitly optimized for speed.
Table 21: Wall-clock runtimes (seconds) required to generate 1 0001\,000 sequences from each model, at the largest batch size which fit in memory.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $16.15$ $20.77$ $69.14$ $22.62$ $1.85$ $6.55$ $9.60$
NHP $0.99$ $6.13$ $11.94$ $1.19$ $0.83$ $0.74$ $0.89$
Diffusion $43.77$ $99.82$ $262.26$ $78.77$ $43.41$ $29.82$ $32.85$
Add-and-Thin $149.09$ $261.53$ $292.51$ $42.60$ $16.71$ $34.73$ $34.87$
EventFlow (1 NFE) $35$ $739$ $885$ 88 $42$ 66 $10$
EventFlow (25 NFEs) $769$ $17889$ $19624$ $187$ $960$ $142$ $233$